# Constraining the Primoridal Black Hole Abundance with Space-Based Detectors

### v1

@@ Acknowledgement
We thank Sachiko Kuroyanagi for early collaboration, and thank Jaume Garriga and Jia-ning Wang for helpful comments and discussions. We also thank Huai-Ke Guo and Jian-dong Zhang for revising our discussion on the recent progress of Taiji and TianQin, respectively. S.P. thanks Chenggang Shao and the Center for Gravitational Experiments of Huazhong University of Science and Technology for the hospitality during his visit. This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, and JSPS KAKENHI grant No. 24K00624. S.P. is also supported by the National Natural Science Foundation of China (NSFC) Grants Nos. 12475066 and 12447101. A.W. is also supported by the DFG under the Emmy-Noether program project number 49659236 and the UCAS Joint PhD Training Program. Z.Z. is also supported by NSFC Grant Nos. 12275004 and 12547127.

### v2

@@ Acknowledgements
We thank Sachiko Kuroyanagi for early collaboration, and thank Jaume Garriga and Jia-ning Wang for helpful comments and discussions. We also thank Huai-Ke Guo and Jian-dong Zhang for revising our discussion on the recent progress of Taiji and TianQin, respectively. S.P. thanks Chenggang Shao and the Center for Gravitational Experiments of Huazhong University of Science and Technology for the hospitality during his visit, and thanks the APCTP workshop “New Perspectives on Cosmology 2026” (APCTP-2026-F02). This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, and JSPS KAKENHI grant No. 24K00624. S.P. is also supported by the National Natural Science Foundation of China (NSFC) Grant Nos. 12475066 and 12447101. A.W. is also supported by the DFG under the Emmy-Noether program project number 49659236 and the UCAS Joint PhD Training Program. Z.Z. is also supported by NSFC Grant Nos. 12275004 and 12547127.